### v1

# The Fourth Conference on System and Service Quality - QualITA, June 25-26, 2025, Catania, Italy

Abstract
As hybrid quantum-classical models gain traction in machine learning, there is a growing need for tools that assess their effectiveness beyond raw accuracy. We present QMetric, a Python package offering a suite of interpretable metrics to evaluate quantum circuit expressibility, feature representations, and training dynamics. QMetric quantifies key aspects such as circuit fidelity, entanglement entropy, barren plateau risk, and training stability. The package integrates with Qiskit and PyTorch, and is demonstrated via a case study on binary MNIST classification comparing classical and quantum-enhanced models. Code, plots, and a reproducible environment are available on GitLab.
keywords:
[…]
1 Introduction
Hybrid quantum-classical neural networks (QNNs) are playing a central role in the development of algorithms for near-term quantum devices. By embedding parameterized quantum circuits within classical training loops, these architectures aim to leverage quantum resources such as entanglement and superposition while maintaining trainability through well-established classical optimizers. This hybrid structure has enabled a broad spectrum of quantum machine learning (QML) models to flourish across domains including classification, generative modeling, optimization, and quantum chemistry.
Importantly, many canonical variational algorithms—originally developed for quantum simulation—can be reframed as learning architectures. The Variational Quantum Eigensolver (VQE) [1, 2] exemplifies this duality: a parameterized quantum ansatz is trained to minimize an energy objective, just like a neural network minimizing a loss function. Over time, VQE has evolved into a family of learning-based formulations, including State-Averaged Orbital-Optimized VQE [3, 4], ADAPT-VQE [5], and Subspace-Search VQE [6], each introducing novel strategies for parameterization, target state selection, and optimization flow.
Beyond simulation, hybrid QNNs are widely applied in supervised learning, typically in classification or regression tasks[7]. Here, quantum circuits are used to encode classical data (via feature maps), process it through a variational ansatz, and output probabilities or decision boundaries. Such architectures are used in Quantum Neural Networks [8], Quantum Support Vector Machines [9], and more recent paradigms like Quantum Kitchen Sinks [10] and Quantum Feature Spaces [11]. In unsupervised learning, models such as Quantum Circuit Born Machines [12], and Quantum Autoencoders [13] extend the reach of QML into generative and latent-variable modeling.
Despite the growing variety and sophistication of QML models, there remains a lack of principled, interpretable, and reproducible tools for evaluating their behavior. Traditional ML diagnostics—accuracy, F1-score, or validation loss—do not capture key quantum characteristics such as circuit expressibility, entanglement structure, barren plateaus, or the sensitivity of quantum feature maps. Without such metrics, model design becomes largely heuristic and comparisons between quantum and classical architectures are often inconclusive or misleading.
To bridge this gap, we introduce QMetric, a modular and extensible Python framework for evaluating hybrid quantum-classical models. QMetric computes interpretable scalar metrics across three complementary dimensions: (i) the structure and expressiveness of quantum circuits; (ii) the geometry and compression of quantum feature spaces; and (iii) the stability, efficiency, and gradient flow during training. These tools allow researchers to diagnose bottlenecks, compare architectures, and validate empirical claims beyond raw accuracy.
Our package integrates with Qiskit  [^1] https://www.ibm.com/quantum/qiskit and PyTorch  [^2] https://pytorch.org/ and is demonstrated through a binary classification example on MNIST digits, comparing a classical neural network to a hybrid QNN. All code, plots, and environment files are publicly available for reproducibility and further experimentation.
2 Metric Categories
QMetric organizes its metrics into three complementary categories—quantum circuit behavior, quantum feature space, and training dynamics—that together provide a comprehensive profile of a hybrid model’s expressiveness, learnability, and robustness. These categories are summarised in Tab. 1.
2.1 Quantum Circuit Metrics
As the computational core of hybrid models, quantum circuits influence representational capacity and noise resilience. QMetric evaluates circuit quality through metrics such as Quantum Circuit Expressibility (QCE), which measures the diversity of quantum states produced under random parameters, and Quantum Circuit Fidelity (QCF), estimating robustness to noise via state overlap.
To characterize circuit structure, the Quantum Locality Ratio (QLR) captures the balance between local and entangling gates. Entanglement-based metrics include Effective Entanglement Entropy (EEE) and Quantum Mutual Information (QMI), which quantify intra-circuit quantum correlations. These metrics are useful when tuning ansätze for VQE, QAOA, or classification tasks, where poor expressibility or excessive entanglement can hinder learning.
2.2 Quantum Feature Space Metrics
When encoding classical data into Hilbert space, the geometry of the resulting feature space directly affects model performance. QMetric provides the Feature Map Compression Ratio (FMCR), assessing how efficiently classical data are compressed via PCA, and the Effective Dimension (EDQFS), which reflects variance spread in the quantum feature space.
The Quantum Layer Activation Diversity (QLAD) and Quantum Output Sensitivity (QOS) evaluate output variability and robustness to perturbations. Low QLAD and high QOS signal collapsed or brittle encodings. These metrics are critical in PQC-based classifiers, quantum kernel methods, and other models relying on quantum feature geometry.
Table 1: Summary of QMetric Metrics
Category Metric Purpose
Quantum Circuit QCE, QCF Expressibility and noise robustness of circuits.
QLR Balance of local vs. entangling gates.
EEE, QMI Entanglement and intra-circuit correlation.
Feature Space FMCR, EDQFS Data compression and feature spread
QLAD Diversity of quantum activations.
QOS Sensitivity to small input changes.
Training Dynamics TSI, TEI Stability and efficiency of convergence.
QGN, BPI Gradient health and barren plateau risk.
RQLSI, r-QTEI Relative diagnostics vs classical models.
2.3 Training Dynamics
QMetric also tracks training behavior using the Training Stability Index (TSI), which compares variability in training and validation losses, and the Training Efficiency Index (TEI), which measures epochs needed to reach a target accuracy relative to model size.
Quantum-specific diagnostics include the Quantum Gradient Norm (QGN) and Barren Plateau Indicator (BPI), both of which expose vanishing gradients linked to deep or poorly initialized circuits. To compare hybrid and classical models, relative metrics such as RQLSI and r-QTEI quantify differences in training efficiency and stability under aligned conditions. Together, these metrics support targeted diagnosis of underperformance, guide ansatz design, and enable meaningful evaluation across model types.
3 Case Study: Hybrid vs Classical on MNIST
To demonstrate QMetric’s diagnostic value, we applied it to a binary classification task on MNIST digits 0 vs 1, comparing a classical feedforward network to a hybrid quantum-classical model using Qiskit’s ZZFeatureMap and RealAmplitudes circuits. Inputs were reduced to 3 dimensions via PCA to match the quantum input size, and both models were trained for 30 epochs using Adam. The classical model reached 99.8% validation accuracy by epoch 5 and stabilized at 99.6%, while the hybrid model plateaued at 69.6% and never surpassed 90%. Despite this gap, the hybrid model showed smoother convergence with a much lower Training Stability Index (TSI = 0.0023 vs. 0.7147), though its Training Efficiency Index (TEI) was infinite due to non-convergence.
QMetric revealed that the hybrid model’s underperformance stemmed not from poor circuit design but from encoding limitations. Quantum Output Sensitivity (QOS = 6.74) was high and Activation Diversity (QLAD = 0.00) was absent, indicating a collapsed feature map insensitive to input variation. Compression (FMCR = 3.0) and effective feature dimension (EDQFS = 1.0) were also low. In contrast, architectural metrics were strong: the circuit showed high expressibility (QCE = 0.94), perfect simulated fidelity (QCF = 1.00), and moderate entanglement (EEE = 0.50) and mutual information (QMI = 1.00). These results highlight QMetric’s ability to pinpoint specific failure modes.
4 Outlook
QMetric provides a structured approach for diagnosing hybrid quantum-classical models beyond conventional performance metrics. It highlights key aspects such as training behavior, encoding robustness, and circuit design quality. Future developments will include migration to Qiskit’s Estimator v2, support for additional platforms like PennyLane, and expanded metric coverage for multi-class tasks and generative models.
5 Availability
All source code, examples, metric definitions, and plotting utilities are available at: https://gitlab.com/illesova.silvie.scholar/qmetric
The repository includes a Conda environment file for reproducibility.
6 Acknowledgments
Finanziato dall’Unione Europea - Next Generation EU, Missione 4 Componente 1 CUP D53D23017570001 e E53D23016370001 e dal Governo italiano (Ministero dell’Università e della Ricerca, PRIN 2022 PNRR) – cod. P2022SELA7: “RECHARGE: monitoRing, tEsting, and CHaracterization of performAnce Regressions” – Decreto Direttoriale n. 1205 del 28/7/2023. TR gratefully acknowledges the funding support by program "Excellence initiative—research university" for the AGH University in Krakow as well as the ARTIQ project: UMO-2021/01/2/ST6/00004 and ARTIQ/0004/2021.
Declaration on Generative AI
During the preparation of this work, the authors used OpenAI ChatGPT (GPT-4) in order to: assist with text formatting, clarify structure, and edit grammar. After using these tools, the author reviewed and edited the content as needed and takes full responsibility for the publication’s content.